OpenStudy (anonymous):

how to take the derivative of the sum of an equation? i will award a medal if you help me get the right answer!

OpenStudy (anonymous):

y is a function y = y(x). C = constant, the derivative(y') of a constant is 0. The formula for the derivative is: y= x^n = > y' = nx^n-1

OpenStudy (anonymous):

\[d/dx \sum_{o}^{\inf} [(x^(2n)))/n!]\]

OpenStudy (anonymous):

x is raised to the power 2n

OpenStudy (anonymous):

\[d/dx \sum_{n=o}^{\inf} [(x^(2n)))/n!]\]